how to find the y-intercept, slope and direction of the line of No. 1.) y=2x No. 2.) y=-2x+2 No. 3) y=1/2x-3 No. 4.)x-3y=5 And the last one No. 5.) 2x+4y=6 PLEASE ANSWER MY QUESTION. THANK YOU VERY MUCH. IT WILL BE VERY MUCH HELPFUL IF YOU ANSWER IT BECAUSE I CAN'T UNDERSTAND IT.
Answer by mananth(16946)

1.) y=2x
compare with y= mx +c
m= slope = 2
y intercept = c= 0
since the slope is positive the line rises from the left to the right.


No. 2.) y=-2x+2
m=-2 = slope
y intercept .
Plug x = 0
y= 2
so y intercept = 2
the line passes through the y axis at 2 and goes up from 4th quadrant to 2nd. Quadrant

No. 3) y=(1/2)x-3
slope = 1/2
y intercept =-3
rises up from III quadrant to I quadrant cutting the y axis at -3
No. 4.)x-3y=5
3y=x-6
y=(1/3)x-2
slope = 1/3
y intercept = -2
rises up from III quadrant to I quadrant cutting the y axis at -2


And the last one No. 5.) 2x+4y=6